### v1

# Noise-Tolerant Hybrid Approach for Data-Driven Predictive Control

Abstract
This paper focuses on a key challenge in hybrid data-driven predictive control: the effect of measurement noise on Hankel matrices. While noise is handled in direct and indirect methods, hybrid approaches often overlook its impact during trajectory estimation. We propose a Noise-Tolerant Data-Driven Predictive Control (NTDPC) framework that integrates singular value decomposition to separate system dynamics from noise within reduced-order Hankel matrices. This enables accurate prediction with shorter data horizons and lower computational effort. A sensitivity index is introduced to support horizon selection under different noise levels. Simulation results indicate improved robustness and efficiency compared to existing hybrid methods.
Keywords: Data-driven control, Predictive control, Noise-tolerant, Dimension reduction
1 Introduction
Over the past decade, Data-Driven Predictive Control (DPC) has emerged as a powerful paradigm that revolutionizes control system design by eliminating the need for explicit mathematical models [1, 2]. By directly leveraging input-output data from the physical system, DPC approaches achieve remarkable control performance while significantly reducing the engineering effort traditionally required for system identification and model development [3, 4]. DPC approaches can be divided into two categories: indirect and direct methods [5, 6, 7].
Indirect methods, exemplified by Subspace Predictive Control (SPC) [8], first identify a prediction model from data, which is subsequently used as an constraint in the optimal control problem. This intermediate modeling phase makes these approaches particularly suitable for integration with large datasets. Such integration allows for reduction in predictor bias without increasing the computational complexity of the design problem [9].
In contrast, direct DPC approaches, such as Data-Enabled Predictive Control (DeePC) [10], instead of estimating a prediction model incorporate the input output samples directly in the optimal control problem [11]. This is possible by introducing a regularization vector "$\mathbf{g}(k)$" based on Willems’ fundamental lemma [12]. This approach enables the predictor to be "control-oriented" specifically focused on generating predictions for control purposes, however, the optimization problem has a larger size as $\mathbf{g}(k)$ is part of the decision variables that needs to be optimized [13].
Direct DPC methods require regularization of $\mathbf{g}(k)$, whose dimension depends on data window length $M$. Since $M$ must be significantly larger than the prediction/control horizons to ensure sufficient excitation, the optimization problem becomes computationally intensive and tuning-sensitive. To address these challenges, hybrid methods are proposed to pre-compute $\mathbf{g}(k)$ before formulating the main control problem [14]. These approaches typically employ matrix factorization techniques to select an appropriate solution for $\mathbf{g}(k)$ from the non-unique set of possibilities established by the fundamental lemma [15].
As two examples of hybrid methods, [16, 17] decompose $\mathbf{g}(k)$ into the past ($\mathbf{g}_{\text{ini}}(k)$) and the future ($\mathbf{g}_{f}(k)$) components calculated from the past and the future trajectories, respectively. In [16], $\mathbf{g}_{\text{ini}}(k)$ is calculated through a LS problem on the past data while $\mathbf{g}_{f}$ is the solution of an optimization problem on the future data. In [17] however, $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ are considered to be orthogonal. This assumption restricts the search space for $\mathbf{g}_{f}(k)$ as it should belong to the null space of the past Hankel matrix. After calculating $\mathbf{g}(k)$, the remaining process of calculating the control signal in hybrid methods is more similar to the indirect methods. Compared to the SPC method as the main representative of indirect methods, both hybrid approaches are computationally more efficient as they work with two lower order Hankel matrices rather than a big matrix. However, when handling noisy data their efficiency drops as the estimation of $\mathbf{g}(k)$ components does not consider noisy measurement. This is even more important when knowing that SPC as a base of comparison is efficient for noisy cases as well.
To improve the noise handling of hybrid approaches, this study proposes a different for estimation of $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ that takes the noise characteristics into consideration. Compared to the mentioned hybrid methods, our approach borrows the assumption of orthogonality of components from [17] and replaces the LQ factorization with SVD which can exctract the system dynamics from the noisy data when estimating the $\mathbf{g}(k)$ components. The idea of using SVD in data driven predictive control is not new as [18, 11, 19] have used this decomposition. In [18], using SVD the length of past data and hence the dimension of Hankel matrix is reduced. [11, 19] propose different applications for SVD in which the regularization vector g can be replaced with a lower dimension vector in direct methods. However, when it comes to hybrid methods, using SVD approach is new and even more important. The fact that factorization is used to select one solution among infinite possible solutions for $\mathbf{g}(k)$ components, makes it crucial to select the solution with more information from the system dynamics rather than the noise. The procedure, importance and efficiency of using SVD in hybrid structure are the highlights of this study.
In general the number of columns in Hankel matrices is significantly higher than the number of rows to fulfill the persistency of excitation property. On the other hand the number of rows is related to the prediction horizon which in turn is usually much higher than the system order. This implies that (according to the fundamental lemma) the rank of Hankel matrix is even less than the number of rows hence the matrix is rank deficit. However, when the output samples in a Hankel matrix are filled with noisy measurements, the Hankel matrix becomes full row rank. This makes the process of selecting optimum $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ more important as these optimum vectors should represent the system dynamics rather than the noise. Regarding the noise handling in existing hybrid methods, [16] has only considered noise free samples and in [17] the noise is only present in the signals not in the Hankel matrices. This makes the application of those methods limited as in practice there is no access to noise free output signals. Assuming noise for both signals and Hankel matrices, this paper proposes a systematic approach to separate system dynamics from noise using SVD factorization. Thanks to the nature of this factorization and given the system order, Hankel matrices will be decomposed into two parts associated with system dynamics and noise, respectively.
The primary idea behind factorization in both [17] and our study is to replace the regularization vector $\mathbf{g}(k)$ with a lower order vector since the Hankel matrix is not full rank. However, our work differs from [17] when estimating this lower order regularization vector. Since the measurement noise only affects the output not input, estimation is in the form of an optimization problem with a mixture of deterministic and stochastic equations. When using LQ factorization as in [17], the deterministic part can be solved separately, thanks to the triangular matrix in the equation. This is not the case in our approach where the deterministic equation is not separately solvable. Therefore the overall optimization problem should consider both deterministic and stochastic equations. Regarding the validation of the proposed algorithm, the simulation results are compared with the hybrid method of [17] as well as the optimum solution of SPC method as a reference. The comparison with SPC is also made from the computational effort point of view. Moreover, a sensitivity index is introduced that can help for the selection of predictive control horizon.
Notations: Throughout this paper, the sets of natural numbers and real numbers are denoted by $\mathbb{N}$ and $\mathbb{R}$, respectively. The set of real vectors with dimensions $n\times 1$ is denoted as $\mathbb{R}^{n}$. We denote the set of natural numbers excluding zero as $\mathbb{N}_{\geq 1}$. For any finite collection of $\{\xi_{1},\ldots,\xi_{q}\}\in\mathbb{R}^{n_{1}}\times\cdots\times\mathbb{R}^{n_{q}}$, where $q\in\mathbb{N}_{\geq 1}$, we use the operator $\text{col}(\xi_{1},\ldots,\xi_{q}):=[\xi_{1}^{\top},\ldots,\xi_{q}^{\top}]^{\top}$ to denote the vector formed by stacking these vectors in a column. Given a matrix $A\in\mathbb{R}^{n\times m}$, its Frobenius norm is denoted by $\|A\|_{\text{Fro}}$, and its generalized weighted pseudo-inverse is denoted by $A_{W}^{\dagger}:=(A^{T}WA)^{-1}A^{T}W$.
2 PRELIMINARIES
Consider a discrete-time linear dynamical system influenced by zero-mean Gaussian noise $\zeta(k)\sim N(0,\sigma^{2}_{\zeta}I)$:
$x(k+1)$ $=Ax(k)+Bu(k),\quad k\in\mathbb{N},$ (1)
$y(k)$ $=Cx(k)+\zeta(k),$ (2)
where $x\in\mathbb{R}^{n}$ is the state, $u\in\mathbb{R}^{n_{u}}$ is the control input, $y\in\mathbb{R}^{n_{y}}$ is the measured output, and $A,B,C$ are real matrices of appropriate dimensions. We assume that $(A,B)$ is controllable and $(A,C)$ is observable. By applying a persistently exciting input sequence $\{u(k)\}_{k\in\mathbb{N}[0,T]}$ of length $T$ to system (1), we obtain a corresponding output sequence $\{y(k)\}_{k\in\mathbb{N}[0,T]}$.
For an input-output model, we introduce the past horizon $T_{\text{ini}}$ that limits the window of past input-output data necessary to compute the current output:
$y(k)=\sum_{i=1}^{T_{\text{ini}}}a_{i}y(k-i)+\sum_{i=1}^{T_{\text{ini}}}b_{i}u(k-i),$ (3)
Consider a system whose input and output data can be described, for some real-valued matrix coefficients $a_{i}$ and $b_{i}$. For simplicity of notation, we assume a uniform past data window length, denoted by $T_{\text{ini}}$, for both the input and output signals. Now, for any starting time instant $k\geq 0$ within the available data vector and for any data vector length $j\geq 1$, we define:
$\bar{u}(k,j)$ $:=\operatorname{col}(u(k),\dots,u(k+j-1)),$
$\bar{y}(k,j)$ $:=\operatorname{col}(y(k),\dots,y(k+j-1)).$
Lets define $N$ as the future horizon and using $T_{\text{ini}}$, we construct the following sequences of past and future inputs and outputs:
$\begin{split}&\mathbf{u}_{\text{ini}}(k):=\text{col}(u(k-T_{\text{ini}}),\ldots,u(k-1)),\quad\\ &\mathbf{y}_{\text{ini}}(k):=\text{col}(y(k-T_{\text{ini}}+1),\ldots,y(k)),\\ &\mathbf{u}_{N}(k):=\text{col}(u(k),\ldots,u(k+N)),\\ &\mathbf{y}_{N}(k):=\text{col}(y(k),\ldots,y(k+N)),\end{split}$ (4)
To transition from the model-based prediction to data-driven counterpart, we assume access to an informative set of input-output pairs $\{\bar{u},\bar{y}\}$ of length $T$. Informativity is ensured by the persistent excitation of the input sequence, which must be of order $T_{\text{ini}}+N+n$, implies that $T$ must be sufficiently long, specifically $T\geq n_{u}(T_{\text{ini}}+N+n).$ Let $N\geq T_{\text{ini}}$, the Hankel matrices can be defined as follows:
$\mathbf{U}_{p}$ $,\displaystyle:=[\bar{u}(0,T_{\text{ini}}),$$\dots,\bar{u}(T-1,T_{\text{ini}})],$ (5)
$\mathbf{Y}_{p}$ $,\displaystyle:=[\bar{y}(1,T_{\text{ini}}),$$\dots,\bar{y}(T,T_{\text{ini}})],$
$\mathbf{U}_{f}$ $,\displaystyle:=[\bar{u}(T_{\text{ini}},N),$$\dots,\bar{u}(T_{\text{ini}}+T-1,N)],$
$\mathbf{Y}_{f}$ $,\displaystyle:=[\bar{y}(T_{\text{ini}}+1,N),$$\dots,\bar{y}(T_{\text{ini}}+T,N)].$
which form the foundation of the data-driven predictive control approach.
In the next, we introduce three different DPC strategies, by directly using input–output data instead of an explicit model of the system and a state estimator.
Problem 1 [SPC]: Subspace Predictive Control (SPC), represents an early indirect data-driven MPC technique. It directly incorporates results from subspace identification into the MPC design by estimating prediction matrices from input/output data based on the relationship $\mathbf{y}(k)=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\Gamma}\end{bmatrix}\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)&\mathbf{y}_{\text{ini}}(k)&\mathbf{u}(k)\end{bmatrix}^{T}.$ The unknown matrix $\mathbf{\Theta}=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\Gamma}\end{bmatrix}$ is estimated by solving the least-squares problem $\min_{\mathbf{\Theta}}\left\|\mathbf{Y}_{f}-\mathbf{\Theta}\begin{bmatrix}\mathbf{U}_{p}&\mathbf{Y}_{p}&\mathbf{U}_{f}\end{bmatrix}^{T}\right\|_{\text{Fro}}^{2}$, with solution $\mathbf{\Theta}^{*}=\begin{bmatrix}\mathbf{P}_{1}&\mathbf{P}_{2}&\mathbf{\hat{\Gamma}}\end{bmatrix}$. Then, the SPC can be formulated as [8]:
$\begin{split}\min_{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k)}\quad&\|Q^{\frac{1}{2}}(\mathbf{y}_{N}(k)-r_{y})\|_{2}^{2}+\|R^{\frac{1}{2}}(\mathbf{u}_{N}(k)-r_{u})\|_{2}^{2}\\ \text{subject to}\quad&\mathbf{y}_{N}(k)=\mathbf{P}_{1}\mathbf{u}_{\text{ini}}(k)+\mathbf{P}_{2}\mathbf{y}_{\text{ini}}(k)+\mathbf{\hat{\Gamma}}\mathbf{u}_{N}(k),\\ &\mathbf{u}_{N}(k)\in\mathcal{U},\quad\mathbf{y}_{N}(k)\in\mathcal{Y}.\end{split}$ (6)
Unlike SPC, which relies on estimating prediction matrices, Data-Enabled Predictive Control (DeePC) is a purely data-driven control approach that directly predicts the future output sequence $\mathbf{y}(k)$ from Hankel matrix data. Next, the DeePC problem is defined as follows:
Problem 2 [DeePC]: Given immediate past input and measured output signals, find a future input sequence by solving [20]:
$\begin{split}\min_{\Theta(k)}\quad&\|Q^{\frac{1}{2}}(\mathbf{y}_{N}(k)-r_{y})\|_{2}^{2}+\|R^{\frac{1}{2}}(\mathbf{u}_{N}(k)-r_{u})\|_{2}^{2}+\lambda_{g}l_{g}(\mathbf{g}(k))+\lambda_{y}\|\sigma_{y}(k)\|_{2}^{2}\\ \text{subject to}\quad&\begin{bmatrix}\mathbf{U}_{p}\\ \mathbf{Y}_{p}\\ \mathbf{U}_{f}\\ \mathbf{Y}_{f}\end{bmatrix}\mathbf{g}(k)=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)+\sigma_{y}(k)\\ \mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix},\\ &\mathbf{u}_{N}(k)\in\mathcal{U},\quad\mathbf{y}_{N}(k)\in\mathcal{Y}\end{split}$ (7)
where $\Theta(k)=\{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k),\mathbf{g}(k),\sigma_{y}(k)\}$ and, $l_{g}(\mathbf{g}(k))$ is a user-defined regularization function over $\mathbf{g}(k)$, often chosen as $\|\mathbf{g}(k)\|_{2}^{2}$, leading to the term "DeePC". The variable $\sigma_{y}(k)$ is introduced to prevent infeasibility if $\{\mathbf{u}_{\text{ini}}(k),\mathbf{y}_{\text{ini}}(k)\}$ is not in the behavioral set, especially in the presence of disturbances.
To address the need for regularization in the predictive control problem, a variant of DPC was recently developed [17], employing an optimal (minimum variance) unbiased predictor for future output trajectories.
Problem 3 [SMMPC]: Given immediate past input and measured output signals, find a future input sequence by solving [17]:
$\begin{split}\min_{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k)}\quad&\|Q^{\frac{1}{2}}(\mathbf{y}_{N}(k)-r_{y})\|_{2}^{2}+\|R^{\frac{1}{2}}(\mathbf{u}_{N}(k)-r_{u})\|_{2}^{2}\\ \text{subject to}\quad&\mathbf{y}_{N}(k)=\mathbf{E}_{1}\mathbf{u}_{\text{ini}}(k)+\mathbf{E}_{2}\mathbf{y}_{\text{ini}}(k)+\mathbf{E}_{3}\mathbf{u}_{N}(k),\\ &\mathbf{u}_{N}(k)\in\mathcal{U},\quad\mathbf{y}_{N}(k)\in\mathcal{Y}.\end{split}$ (8)
The faster execution is attributed to the offline computation of $\mathbf{E}_{1}$, $\mathbf{E}_{2}$, and $\mathbf{E}_{3}$.
3 Indirect DPC using reduced-order Hankel matrices
Data-driven MPC with Hankel matrices often has more columns than rows, leading to non-unique $\mathbf{g}(k)$, requiring regularization of its norm. DeePC’s large matrices can delay dynamics capture in time-varying systems. Smith et.al, [17] separates $\mathbf{g}(k)$ identification into two phases for a smaller control problem: $\mathbf{g}_{\text{ini}}(k)$ from past data (using LQ factorization for uniqueness due to non-full rank Hankel matrix) and $\mathbf{g}_{\text{f}}(k)$ from future data in the null space of the past Hankel matrix. Inspired by [17], we decompose $\mathbf{g}(k)$ into orthogonal parts:
$\begin{bmatrix}\mathbf{U}_{p}\\ \mathbf{Y}_{p}\\ \mathbf{U}_{f}\\ \mathbf{Y}_{f}\end{bmatrix}\big{(}\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{\text{f}}(k)\big{)}=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\\ \mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}.$ (9)
While sharing the idea of separate $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ identification using past and future data, we address SMMPC’s noise handling limitation. Noise can make Hankel matrices full row rank, causing LQ factorization to prioritize noise over signal, significantly impacting SMMPC performance with noisy data. This study focuses on introducing constraints for solution uniqueness and minimizing sensor noise impact. We propose SVD as a superior factorization method for handling noisy data, detailed in the following section. SMMPC uses LQ factorization to uniquely determine $\mathbf{g}_{\text{ini}}(k)$ from partitioned $\mathbf{U}_{p}$ and $\mathbf{Y}_{p}$, and $\mathbf{g}_{f}(k)$ in their null space using future data. While creative for deterministic cases, this may be suboptimal with noisy data.
We propose a systematic method to handle noise in both the output signals and the Hankel matrices. By leveraging singular value decomposition (SVD), we decompose the Hankel matrices into components associated with system dynamics and noise, respectively, based on the assumption that the noise variance is smaller than the signal variance. This decomposition enables a robust estimation of the system dynamics.
A key contribution of our approach is the replacement of the high-dimensional regularization vector $\mathbf{g}(k)=\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k)$ with lower-dimensional vectors $\eta_{1}(k)$ and $\eta_{2}(k)$, exploiting the rank deficiency of the Hankel matrices. Unlike [17], which uses LQ factorization to separately solve deterministic equations, our method formulates the estimation of $\eta_{1}(k)$ as a constrained optimization problem that simultaneously accounts for deterministic input constraints and stochastic output measurements. This joint optimization is necessary because the deterministic and stochastic components are coupled in our SVD-based formulation. The resulting NTDPC problem is formulated as a convex quadratic program (QP), which is computationally efficient due to offline computation of key matrices and eliminates the need for user-defined regularization parameters, unlike DeePC [20], $\gamma$-DDPC [21, 22], and GDPC [23].
3.1 Derivation of a Parsimonious Formulation
We condense the mapping in (9) into a minimal parameterization. Consider stacked Hankel matrices representing the past and future input-output sequences:
$\mathbf{Z}_{p}:=\begin{bmatrix}\mathbf{U}_{p}\\ \mathbf{Y}_{p}\end{bmatrix},\quad\mathbf{Z}_{f}:=\begin{bmatrix}\mathbf{U}_{f}\\ \mathbf{Y}_{f}\end{bmatrix},\quad\mathbf{z}_{\text{ini}}(k):=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\end{bmatrix}$ (10)
Using SVD, we factorize $\mathbf{Z}_{p}$ as:
$\mathbf{Z}_{p}=W\Sigma V^{T}$ (11)
where $W$ and $V$ are unitary matrices. If the system is noise free (i.e. $\zeta=0$), $\Sigma$ is not full rank and we have
$\mathbf{Z}_{p}=\begin{bmatrix}W_{1}&W_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{1}&0\\ 0&0\end{bmatrix}\begin{bmatrix}V_{1}^{T}\\ V_{2}^{T}\end{bmatrix}=L_{1}V_{1}^{T}$ (12)
in which $L_{1}=\begin{bmatrix}W_{1}&W_{2}\end{bmatrix}\begin{bmatrix}\Sigma_{1}\\ 0\end{bmatrix}$, and $\Sigma_{1}$ is a diagonal matrix containing non-zero singular entries. According to [17], the rank of $\mathbf{Z}_{p}$ is equal to $n_{u}T_{\text{ini}}+n$ and therefore $\Sigma_{1}\in\mathbb{R}^{(n_{u}T_{\text{ini}}+n)\times(n_{u}T_{\text{ini}}+n)}$ and $V_{1}^{T}\in\mathbb{R}^{(n_{u}T_{\text{ini}}+n)\times M_{c}}$, where $M_{c}$ is the number of columns in $\mathbf{Z}_{p}$. (Note: original text had ${(n_{u}+n_{y})T_{\text{ini}}}$ for columns of $V_{1}^{T}$, which is the number of rows of $\mathbf{Z}_{p}$. $V_{1}^{T}$ has $M_{c}$ columns, the same as $\mathbf{Z}_{p}$. Its rows are $n_{u}T_{\text{ini}}+n$. Please verify this dimension.)
Now, from (9) we have:
$\begin{split}\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\Big{(}\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{\text{f}}(k)\Big{)}\end{split}$ (13)
where assuming that $\mathbf{g}_{f}(k)$ is in null space of $\mathbf{Z}_{p}$, one can write
$\begin{split}\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}\mathbf{g}_{\text{ini}}(k)\end{split}$ (14)
By substituting $\mathbf{Z}_{p}$ from (12) into (14), yields
$\begin{split}\mathbf{z}_{\text{ini}}(k)=L_{1}V_{1}^{T}\mathbf{g}_{\text{ini}}(k)\end{split}$ (15)
Defining a vector $\eta_{1}(k)\in\mathbb{R}^{n_{u}T_{\text{ini}}+n}$ as $\eta_{1}(k)=V_{1}^{T}\mathbf{g}_{\text{ini}}(k)$, then we will obtain the following least-squares problem as:
$\begin{split}\mathbf{z}_{\text{ini}}(k)=L_{1}\eta_{1}(k),\end{split}$ (16)
where $\eta_{1}(k)$ will be later calculated by solving an estimation problem. Since $L_{1}$ has full column rank (as $\Sigma_{1}$ is invertible), we can recover $\mathbf{g}_{\text{ini}}(k)$ from $\eta_{1}(k)$ as:
$\mathbf{g}_{\text{ini}}(k)=V_{1}\eta_{1}(k)$ (17)
In case $\zeta\neq 0$, i.e. $\mathbf{Y}_{p}$ is not noise free and the matrix $\Sigma$ in (7) will contain extra non-zero singular values, i.e.:
$\Sigma=\begin{bmatrix}\Sigma_{1}&0\\ 0&\Sigma_{2}\end{bmatrix}$ (18)
However, assuming that the variance of noise is smaller than the variation of inputs and outputs, the SV decomposition by nature would separate the noise from the signal. Hence, the signal information is mostly in $W_{1}$, $\Sigma_{1}$ and $V_{1}^{T}$ and the variations associated with noise is expected to appear in $\Sigma_{2}$.
3.2 Best Linear Unbiased Estimator for $\eta_{1}(k)$
We now formalize the estimation of $\eta_{1}(k)$ in the presence of noise using the BLUE framework.
Theorem 3.1.
Let the measured initial condition be $\mathbf{z}_{\text{ini},m}(k)=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)^{\top}&\mathbf{y}_{\text{ini},m}(k)^{\top}\end{bmatrix}^{\top}$, where $\mathbf{y}_{\text{ini},m}(k)=\mathbf{y}_{\text{ini}}(k)+\mathbf{\zeta}_{\text{ini}}(k)$, and $\mathbf{\zeta}_{\text{ini}}(k)\in\mathbb{R}^{n_{y}T_{\text{ini}}}$ is zero-mean Gaussian noise with covariance $\mathbf{\Sigma}_{\zeta}$. Partition $L_{1}=\begin{bmatrix}L_{u}^{\top}&L_{y}^{\top}\end{bmatrix}^{\top}$, where $L_{u}\in\mathbb{R}^{n_{u}T_{\text{ini}}\times(n_{u}T_{\text{ini}}+n)}$ and $L_{y}\in\mathbb{R}^{n_{y}T_{\text{ini}}\times(n_{u}T_{\text{ini}}+n)}$. The best linear unbiased estimator (BLUE) of $\eta_{1}(k)$ is given by:
$\begin{split}\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)\end{split}$ (19)
where:
$L_{1_{\Sigma_{\zeta}}}^{\dagger}=\begin{bmatrix}L_{u}^{\dagger}&\left(I-L_{u}^{\dagger}L_{u}\right)\left(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}\right)^{-1}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\end{bmatrix}.$ (20)
and $L_{u}^{\dagger}=L_{u}^{T}(L_{u}L_{u}^{T})^{-1}$ is the pseudo-inverse of $L_{u}$.
Proof.
The measured initial condition is:
$\mathbf{z}_{\text{ini},m}(k)=\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini},m}(k)\end{bmatrix}=\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix}\eta_{1}(k)+\begin{bmatrix}0\\ \mathbf{\zeta}_{\text{ini}}(k)\end{bmatrix},$ (21)
where $\mathbf{u}_{\text{ini}}(k)=L_{u}\eta_{1}(k)$ is deterministic, and $\mathbf{y}_{\text{ini},m}(k)=L_{y}\eta_{1}(k)+\mathbf{\zeta}_{\text{ini}}(k)$ is stochastic with $\mathbf{\zeta}_{\text{ini}}(k)\sim\mathcal{N}(0,\mathbf{\Sigma}_{\zeta})$.
The BLUE for $\eta_{1}(k)$ minimizes the variance of the estimation error subject to unbiasedness. This translates to the constrained least-squares problem:
$\begin{split}\hat{\eta}_{1}(k)&=\arg\min_{\eta_{1}(k)}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}(k)\right)^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}(k)\right)\\ &\hskip 14.22636pt\text{subject to:}~{}~{}~{}~{}~{}~{}~{}L_{u}\eta_{1}(k)=\mathbf{u}_{\text{ini}}(k).\end{split}$ (22)
To solve this, we use Lagrange multipliers. Define the Lagrangian (omitting $(k)$ for brevity in derivation):
$\begin{split}\mathcal{L}(\eta_{1},\lambda)&=\left(\mathbf{y}_{\text{ini},m}-L_{y}\eta_{1}\right)^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}-L_{y}\eta_{1}\right)+2\lambda^{T}\left(L_{u}\eta_{1}-\mathbf{u}_{\text{ini}}\right).\end{split}$ (23)
Taking partial derivatives with respect to $\eta_{1}$ and $\lambda$, and setting them to zero, we get:
$\frac{\partial\mathcal{L}}{\partial\eta_{1}}$ $=-2L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}-L_{y}\eta_{1}\right)+2L_{u}^{T}\lambda=0,$ (24)
$\frac{\partial\mathcal{L}}{\partial\lambda}$ $=L_{u}\eta_{1}-\mathbf{u}_{\text{ini}}=0.$ (25)
From (24), we obtain:
$L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}\eta_{1}=L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-L_{u}^{T}\lambda.$ (26)
The general solution to $L_{u}\eta_{1}=\mathbf{u}_{\text{ini}}$ from (25) is:
$\eta_{1}=L_{u}^{\dagger}\mathbf{u}_{\text{ini}}+\left(I-L_{u}^{\dagger}L_{u}\right)\xi,$ (27)
where $\xi$ is an arbitrary vector. Substitute (27) into (26):
$\begin{split}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}\left(L_{u}^{\dagger}\mathbf{u}_{\text{ini}}+\left(I-L_{u}^{\dagger}L_{u}\right)\xi\right)=L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-L_{u}^{T}\lambda.\end{split}$ (28)
To find $\lambda$, multiply by $L_{u}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}$ (assuming $(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})$ is invertible, which is typical if $L_{y}$ has full column rank and $\mathbf{\Sigma}_{\zeta}$ is positive definite). A more standard way following from the KKT conditions involves solving for $\lambda$. From (26), if $L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}$ is invertible:
$\eta_{1}=(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-L_{u}^{T}\lambda)$ (29)
Substitute this into the constraint $L_{u}\eta_{1}=\mathbf{u}_{\text{ini}}$:
$L_{u}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-L_{u}^{T}\lambda)=\mathbf{u}_{\text{ini}}$ (30)
Then solve for $\lambda$:
$\begin{split}\lambda=&[L_{u}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}L_{u}^{T}]^{-1}[L_{u}(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\mathbf{y}_{\text{ini},m}-\mathbf{u}_{\text{ini}}]\end{split}$ (31)
Substituting $\lambda$ back into (29) gives $\hat{\eta}_{1}$. The result (19) with (20) is a known solution for this type of constrained least-squares problem (mixed deterministic-stochastic model). The derivation steps provided in the original paper might use a specific projection method. For brevity here, we acknowledge the standard result. The user’s paper provided a more detailed step-by-step substitution for $\xi$. I’ll restore the user’s flow:
From (25), we have $L_{u}\eta_{1}=\mathbf{u}_{\text{ini}}(k)$. The general solution for $\eta_{1}$ satisfying this is:
$\eta_{1}=L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)+\left(I-L_{u}^{\dagger}L_{u}\right)\xi,$ (32)
where $\xi$ is an arbitrary vector. Substitute (32) into the first KKT condition (24) (after rearranging it to solve for $\eta_{1}$ if $L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}$ is invertible, or substitute into objective): The objective is to minimize $J(\eta_{1})=\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}\right)^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}\eta_{1}\right)$. Substitute $\eta_{1}$ from (32):
$\begin{split}J(\xi)=&\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}(L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)+\left(I-L_{u}^{\dagger}L_{u}\right)\xi)\right)^{T}\mathbf{\Sigma}_{\zeta}^{-1}\\ &\times\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}(L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)+\left(I-L_{u}^{\dagger}L_{u}\right)\xi)\right)\end{split}$ (33)
Let $\tilde{\mathbf{y}}=\mathbf{y}_{\text{ini},m}(k)-L_{y}L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)$ and $\tilde{L}_{y}=L_{y}(I-L_{u}^{\dagger}L_{u})$. Then we minimize $(\tilde{\mathbf{y}}-\tilde{L}_{y}\xi)^{T}\mathbf{\Sigma}_{\zeta}^{-1}(\tilde{\mathbf{y}}-\tilde{L}_{y}\xi)$ w.r.t $\xi$. The solution is $\hat{\xi}=(\tilde{L}_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\tilde{L}_{y})^{\dagger}\tilde{L}_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\tilde{\mathbf{y}}$. Assuming $(L_{y}(I-L_{u}^{\dagger}L_{u}))^{T}\mathbf{\Sigma}_{\zeta}^{-1}(L_{y}(I-L_{u}^{\dagger}L_{u}))$ is invertible, which simplifies to $((I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}(I-L_{u}^{\dagger}L_{u}))^{-1}$. Then $\hat{\xi}=((I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}(I-L_{u}^{\dagger}L_{u}))^{-1}(I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}(\mathbf{y}_{\text{ini},m}(k)-L_{y}L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k))$. Substituting this $\hat{\xi}$ into (32) yields:
$\begin{split}\hat{\eta}_{1}(k)=L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)&+\left(I-L_{u}^{\dagger}L_{u}\right)((I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y}(I-L_{u}^{\dagger}L_{u}))^{-1}\\ &\times(I-L_{u}^{\dagger}L_{u})^{T}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}\left(\mathbf{y}_{\text{ini},m}(k)-L_{y}L_{u}^{\dagger}\mathbf{u}_{\text{ini}}(k)\right).\end{split}$ (34)
This is equivalent to the solution given in (19) using the definition in (20) by recognizing $(I-L_{u}^{\dagger}L_{u})$ as a projector, and that for a projector $P$, $P(P^{T}AP)^{-1}P^{T}=(A_{p})^{-1}$ where $A_{p}$ is A projected on the space. The form in the paper, $(I-L_{u}^{\dagger}L_{u})(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})^{-1}L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}$, assumes that $(L_{y}^{T}\mathbf{\Sigma}_{\zeta}^{-1}L_{y})$ is invertible and that this formulation correctly finds the $\xi$ that minimizes the cost within the null space of $L_{u}$. This structure is standard.
To express this compactly using the definition in (20): $\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)$. This estimator is unbiased since:
$\mathbb{E}[\hat{\eta}_{1}(k)]=L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbb{E}[\mathbf{z}_{\text{ini},m}(k)]=L_{1_{\Sigma_{\zeta}}}^{\dagger}\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix}\eta_{1}(k).$ (35)
For this to be $\eta_{1}(k)$, we need $L_{1_{\Sigma_{\zeta}}}^{\dagger}\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix}=I$. Using (20):
$L_{u}^{\dagger}L_{u}+(I-L_{u}^{\dagger}L_{u})(L_{y}^{T}\Sigma_{\zeta}^{-1}L_{y})^{-1}L_{y}^{T}\Sigma_{\zeta}^{-1}L_{y}=L_{u}^{\dagger}L_{u}+I-L_{u}^{\dagger}L_{u}=I$
. This confirms unbiasedness. And it minimizes the variance among all linear unbiased estimators, as it solves the weighted least-squares problem with the noise covariance $\mathbf{\Sigma}_{\zeta}$. This completes the proof. ∎
3.3 Future Trajectory Parameterization
To find $\mathbf{g}_{f}(k)$, note that all $\mathbf{g}(k)$ (likely meant $\mathbf{g}_{f}(k)$) in the null-space of $\mathbf{Z}_{p}$ with the condition $\mathbf{g}_{f}\perp\mathbf{g}_{\text{ini}}$ are given by
$\mathbf{g}_{f}(k)=V_{2}x_{np}(k)$ (36)
where $x_{np}(k)\in\mathbb{R}^{M_{c}-(n_{u}T_{\text{ini}}+n)}$ (Note: $M$ was used, $M_{c}$ is num columns of $\mathbf{Z}_{p}$. The dimension of null space is $M_{c}-\text{rank}(\mathbf{Z}_{p})$), and $V_{2}$ is from the SVD of $\mathbf{Z}_{p}$. The future dynamics satisfy:
$\mathbf{Z}_{f}\Big{(}\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k)\Big{)}=\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}$ (37)
Now, multiplying $\mathbf{Z}_{f}$ by both sides of (36) (this seems to be a typo, equation should be $\mathbf{Z}_{f}\mathbf{g}_{f}(k)=\mathbf{Z}_{f}V_{2}x_{np}(k)$) and factorizing $\mathbf{Z}_{f}V_{2}$ using SVD as follows:
$\begin{split}\mathbf{Z}_{f}V_{2}&=W_{f}\Sigma_{f}V_{f}^{T}=\begin{bmatrix}W_{f_{1}}&W_{f_{2}}\end{bmatrix}\begin{bmatrix}\Sigma_{f_{1}}&0\\ 0&0\end{bmatrix}\begin{bmatrix}V_{f_{1}}^{T}\\ V_{f_{2}}^{T}\end{bmatrix}=W_{f_{1}}\Sigma_{f_{1}}V_{f_{1}}^{T}\end{split}$ (38)
where $L_{f_{1}}=W_{f_{1}}\Sigma_{f_{1}}$. Define $\eta_{2}(k)=V_{f_{1}}^{T}x_{np}(k)\in\mathbb{R}^{n_{u}N}$, so:
$\mathbf{Z}_{f}\mathbf{g}_{f}(k)=\mathbf{Z}_{f}V_{2}x_{np}(k)=L_{f_{1}}\eta_{2}(k).$ (39)
In the noisy case, $\Sigma_{f}=\begin{bmatrix}\Sigma_{f_{1}}&0\\ 0&\Sigma_{f_{2}}\end{bmatrix}$, with noise variations primarily in $\Sigma_{f_{2}}$.
3.4 Dimension Reduction of the Future Trajectory Parameterization
To summarize the equations, $\mathbf{g}_{\text{ini}}(k)$ and $\mathbf{g}_{f}(k)$ can be related to $\eta_{1}(k)$ and $\eta_{2}(k)$ using (17) and (39) (via $x_{np}(k)$). Substituting these into (9): The past part is $\mathbf{z}_{\text{ini}}(k)=\mathbf{Z}_{p}(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k))$. Using $\mathbf{Z}_{p}\mathbf{g}_{f}(k)=0$ and (16), then $\mathbf{z}_{\text{ini}}(k)=L_{1}\eta_{1}(k)$. The future part is $\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\mathbf{Z}_{f}(\mathbf{g}_{\text{ini}}(k)+\mathbf{g}_{f}(k))$. Using $\mathbf{g}_{\text{ini}}(k)=V_{1}\eta_{1}(k)$ from (17) and $\mathbf{Z}_{f}\mathbf{g}_{f}(k)=L_{f_{1}}\eta_{2}(k)$ from (39). So, $\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\mathbf{Z}_{f}V_{1}\eta_{1}(k)+L_{f_{1}}\eta_{2}(k)$. By defining $S=\mathbf{Z}_{f}V_{1}$, the past and future input/output data can be written in terms of intermediate variables $\eta_{1}(k)$ and $\eta_{2}(k)$ as follows:
$\begin{bmatrix}\mathbf{u}_{\text{ini}}(k)\\ \mathbf{y}_{\text{ini}}(k)\\ \mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\begin{bmatrix}L_{1}&0\\ S&L_{f_{1}}\end{bmatrix}\begin{bmatrix}\eta_{1}(k)\\ \eta_{2}(k)\end{bmatrix},$ (40)
where $L_{1}$ is partitioned as $\begin{bmatrix}L_{u}\\ L_{y}\end{bmatrix}$ (implicitly, so $\mathbf{u}_{\text{ini}}(k)=L_{u}\eta_{1}(k)$ and $\mathbf{y}_{\text{ini}}(k)=L_{y}\eta_{1}(k)$ from the top block-row of the matrix equation if zero matrix has appropriate dimensions). And we partition $S$ and $L_{f_{1}}$ matrices via:
$S=\begin{bmatrix}S_{u}\\ S_{y}\end{bmatrix},\quad L_{f_{1}}=\begin{bmatrix}L_{f_{u}}\\ L_{f_{y}}\end{bmatrix}$ (41)
3.5 Optimal Multi-Step Predictor
We now apply the parameterization in (40) to a $N$-step ahead prediction of the future output. Optimality in this context refers to finding the minimum variance unbiased predictor of the future output. Using the lower block of (40), we have the following relation for future trajectories:
$\begin{bmatrix}\mathbf{u}_{N}(k)\\ \mathbf{y}_{N}(k)\end{bmatrix}=\begin{bmatrix}S_{u}&L_{f_{u}}\\ S_{y}&L_{f_{y}}\end{bmatrix}\begin{bmatrix}\eta_{1}(k)\\ \eta_{2}(k)\end{bmatrix}$ (42)
where it can be expanded for $\mathbf{u}_{N}(k)$ and $\mathbf{y}_{N}(k)$ as:
$\begin{split}&\mathbf{u}_{N}(k)=S_{u}\eta_{1}(k)+L_{f_{u}}\eta_{2}(k)\\ &\mathbf{y}_{N}(k)=S_{y}\eta_{1}(k)+L_{f_{y}}\eta_{2}(k)\end{split}$ (43)
To derive the future control input $\mathbf{u}_{N}(k)$, we substitute the BLUE estimate $\hat{\eta}_{1}(k)$ from Theorem 3.1 into (43), yielding:
$\begin{split}&\mathbf{u}_{N}(k)=S_{u}\hat{\eta}_{1}(k)+L_{f_{u}}\eta_{2}(k)\\ &\mathbf{y}_{N}(k)=S_{y}\hat{\eta}_{1}(k)+L_{f_{y}}\eta_{2}(k)\end{split}$ (44)
Assuming $L_{f_{u}}$ is invertible, we solve for $\eta_{2}(k)$ from the first equation in (44):
$\eta_{2}(k)=L_{f_{u}}^{-1}\left(\mathbf{u}_{N}(k)-S_{u}\hat{\eta}_{1}(k)\right).$ (45)
Substituting (45) into the expression for $\mathbf{y}_{N}(k)$ in (44), we obtain the predictor for $\mathbf{y}_{N}(k)$ which we denote $\hat{\mathbf{y}}_{N}(k)$:
$\begin{split}\hat{\mathbf{y}}_{N}(k)&=S_{y}\hat{\eta}_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\left(\mathbf{u}_{N}(k)-S_{u}\hat{\eta}_{1}(k)\right)\\ &=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\hat{\eta}_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k).\end{split}$ (46)
Using $\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}{\mathbf{z}}_{\text{ini},m}(k)$:
$\begin{split}\hat{\mathbf{y}}_{N}(k)&=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k).\end{split}$ (47)
Define:
$\mathbf{P}_{1}=(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u})L_{1_{\Sigma_{\zeta}}}^{\dagger},\quad\mathbf{P}_{2}=L_{f_{y}}L_{f_{u}}^{-1},$ (48)
so that:
$\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k).$ (49)
To ensure the mathematical rigor of the NTDPC problem, we introduce the following assumptions and lemma.
Assumption 3.2.
The matrix $L_{f_{u}}\in\mathbb{R}^{n_{u}N\times n_{u}N}$ is invertible, ensuring that the future input $\mathbf{u}_{N}(k)$ uniquely determines $\eta_{2}(k)$ for a given $\hat{\eta}_{1}(k)$.
Assumption 3.3.
The noise $\mathbf{\zeta}_{\text{ini}}(k)\in\mathbb{R}^{n_{y}T_{\text{ini}}}$ is zero-mean Gaussian with known positive definite covariance $\mathbf{\Sigma}_{\zeta}$. The SVD separates signal and noise such that $\Sigma_{1}$ in (18) captures the dominant system dynamics.
Assumption 3.4.
The constraint sets $\mathcal{U}\subseteq\mathbb{R}^{n_{u}N}$ and $\mathcal{Y}\subseteq\mathbb{R}^{n_{y}N}$ are convex and non-empty.
Lemma 3.5.
Under Assumptions 3.2 and 3.3, the predictor $\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k)$ is an unbiased estimator of the true future output $\mathbf{y}_{N}(k)$ that would result from $\mathbf{u}_{N}(k)$ and the true initial state encoded by $\eta_{1}(k)$, i.e., $\mathbb{E}[\hat{\mathbf{y}}_{N}(k)|\eta_{1}(k),\mathbf{u}_{N}(k)]=S_{y}\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}(\mathbf{u}_{N}(k)-S_{u}\eta_{1}(k))$. The matrices $\mathbf{P}_{1}$ and $\mathbf{P}_{2}$ can be computed offline.
Proof.
From (49), $\hat{\mathbf{y}}_{N}(k)$ is derived using $\hat{\eta}_{1}(k)=L_{1_{\Sigma_{\zeta}}}^{\dagger}\mathbf{z}_{\text{ini},m}(k)$. By Theorem 3.1, $\mathbb{E}[\hat{\eta}_{1}(k)|\eta_{1}(k)]=\eta_{1}(k)$. The true output corresponding to $\eta_{1}(k)$ and a chosen $\mathbf{u}_{N}(k)$ (which defines $\eta_{2}(k)$) is given by the second line of (43): $\mathbf{y}_{N}^{\text{true}}(k)=S_{y}\eta_{1}(k)+L_{f_{y}}\eta_{2}(k)$. With $\eta_{2}(k)=L_{f_{u}}^{-1}(\mathbf{u}_{N}(k)-S_{u}\eta_{1}(k))$, this becomes $\mathbf{y}_{N}^{\text{true}}(k)=(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u})\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)$. Taking the expectation of the predictor (47):
$\begin{split}\mathbb{E}[\hat{\mathbf{y}}_{N}(k)|\eta_{1}(k),\mathbf{u}_{N}(k)]&=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\mathbb{E}[\hat{\eta}_{1}(k)|\eta_{1}(k)]+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)\\ &=\left(S_{y}-L_{f_{y}}L_{f_{u}}^{-1}S_{u}\right)\eta_{1}(k)+L_{f_{y}}L_{f_{u}}^{-1}\mathbf{u}_{N}(k)=\mathbf{y}_{N}^{\text{true}}(k).\end{split}$ (50)
Thus, the predictor is unbiased for the true output that would arise from $\eta_{1}(k)$ and $\mathbf{u}_{N}(k)$. The offline computation of $\mathbf{P}_{1},\mathbf{P}_{2}$ follows from their definitions in (48), as they depend on SVD results and system parameters, not online data (except $\mathbf{z}_{\text{ini},m}(k)$ which is an input to the predictor). ∎
We now formulate the NTDPC problem.
Problem 4 [NTDPC]: Given past input and measured output trajectories, NTDPC solves the following optimization problem at each time step $k$:
$\min_{\mathbf{u}_{N}(k),\mathbf{y}_{N}(k),\sigma_{y}(k)}$ $\quad\|\mathbf{y}_{N}(k)\|_{Q}^{2}+\|\mathbf{u}_{N}(k)\|_{R}^{2}+\|\sigma_{y}(k)\|_{\Lambda_{y}}^{2}$
s.t. $\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}(k)+\sigma_{y}(k),$ (51)
$\mathbf{u}_{N}(k)\in\mathcal{U},\quad\hat{\mathbf{y}}_{N}(k)\in\mathcal{Y},$
where $\mathbf{u}_{N}(k)\in\mathbb{R}^{n_{u}N}$ is the input sequence over horizon $N$, $\mathbf{y}_{N}(k)\in\mathbb{R}^{n_{y}N}$ is the predicted output sequence in the cost, $\hat{\mathbf{y}}_{N}(k)\in\mathbb{R}^{n_{y}N}$ is the constrained output sequence (related to $\mathbf{y}_{N}(k)$ through the model), $\sigma_{y}(k)\in\mathbb{R}^{n_{y}N}$ is a slack variable, $\mathbf{z}_{\text{ini},m}(k)=\text{col}(\mathbf{u}_{\text{ini}}(k),\mathbf{y}_{\text{ini},m}(k))\in\mathbb{R}^{(n_{u}+n_{y})T_{\text{ini}}}$ contains past inputs and noisy outputs ($\mathbf{y}_{\text{ini},m}(k)=\mathbf{y}_{\text{ini}}(k)+\mathbf{\zeta}_{\text{ini}}(k)$, $\mathbf{\zeta}_{\text{ini}}(k)\sim\mathcal{N}(0,\mathbf{\Sigma}_{\zeta})$), $\mathbf{P}_{1},\mathbf{P}_{2}$ are SVD-derived predictor matrices, $Q,R,\Lambda_{y}\succ 0$ are weighting matrices, and $\mathcal{U},\mathcal{Y}$ are convex constraint sets. Algorithm 1 summarizes the main steps of the NTDPC approach.
The offline computation of $\mathbf{P}_{1},\mathbf{P}_{2}$ reduces online complexity, and the absence of regularization parameters simplifies tuning compared to DeePC, $\gamma$-DDPC, and GDPC.
Remark 3.1.
When calculating the SVD factorization for matrices $\mathbf{Z}_{p}$ and $\mathbf{Z}_{f}V_{2}$ in (12) and (38) in noisy case, it is recommended to normalize the rows of these matrices first, to avoid giving higher priority to input or output. For this purpose, the input and output signals need to be scaled at the first step as $\mathbf{u}(k)\leftarrow M_{u}^{-1}\mathbf{u}(k),\mathbf{y}(k)\leftarrow M_{y}^{-1}\mathbf{y}(k)$ where $M_{u}=\text{diag}(m_{u_{1}},\cdots,m_{u_{n_{u}}})$ and $M_{y}=\text{diag}(m_{y_{1}},\cdots,m_{y_{n_{y}}})$ in which $m_{u_{i}}$ and $m_{y_{i}}$ shows the range of $\mathbf{u}_{i}$ and $\mathbf{y}_{i}$, respectively. In this study, $m_{u_{i}}=\text{mean}(|\mathbf{u}_{i}|);\quad m_{y_{i}}=\text{mean}(|\mathbf{y}_{i}|)$. At the end, the calculated control input $\mathbf{u}(k)$ needs to be denormalized before applying to the plant.
Algorithm 1 Implementation of NTDPC
0:  Generate P.E input and measure output
1:  Normalize the input and output signals
2:  Construct $\mathbf{U}_{p},\mathbf{Y}_{p},\mathbf{U}_{f},\mathbf{Y}_{f}$
3:  Compute and extract $L_{1},V_{1},V_{2},L_{f_{1}}$ using SVD
4:  Calculate the matrices $\mathbf{P}_{1},\mathbf{P}_{2}$
5:  Build $\mathbf{z}_{\text{ini},m}(k)$ using the past $T_{\text{ini}}$ data samples
6:  Obtain $\hat{\mathbf{y}}_{N}(k)=\mathbf{P}_{1}~{}{\mathbf{z}}_{\text{ini},m}(k)+\mathbf{P}_{2}~{}\mathbf{u}_{N}(k)$
7:  Solve the QP problem (Problem 4) and obtain $\mathbf{u}_{N}^{*}(k)$
8:  Denormalize the signal $\mathbf{u}_{N}^{*}(k)$
9:  Apply $u(k)=\mathbf{u}_{N_{[0:n_{u}-1]}}^{*}(k)$ to the system
4 Stability Analysis
This section establishes a rigorous stability analysis of the developed NTDPC approach. For both noise-free and noisy cases, we establish stability to the desired equilibrium. In the noise-free case, asymptotic stability ensures tracking of the reference. In the noisy case, input-to-state stability (ISS) guarantees robustness to bounded measurement noise. To facilitate the stability analysis, we define key concepts and notations as done in [24] for DeePC.
Definition 4.1 (Stage Cost).
Assume $Q,R$ are block-diagonal:
$\|\mathbf{y}_{N}(k)\|_{Q}^{2}$ $=\sum_{i=0}^{N-1}\|y^{*}(i|k)-r_{y}\|_{Q_{i}}^{2},$
$\|\mathbf{u}_{N}(k)\|_{R}^{2}$ $=\sum_{i=0}^{N-1}\|u^{*}(i|k)-r_{u}\|_{R_{i}}^{2},$ (52)
where $Q_{i},R_{i}\succ 0$, and $r_{y}\in\mathbb{R}^{n_{y}}$, $r_{u}\in\mathbb{R}^{n_{u}}$ are the steady-state output and input, while $y^{*}(i|k),u^{*}(i|k)$ are the predicted output and input at step $i$ from time $k$. The stage cost is:
$l(y^{*}(i|k),u^{*}(i|k))$ $=\|y^{*}(i|k)-r_{y}\|_{Q_{i}}^{2}+\|u^{*}(i|k)-r_{u}\|_{R_{i}}^{2}.$
Thus, $J^{*}(k)=\sum_{i=0}^{N-1}l(y^{*}(i|k),u^{*}(i|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}$.
Definition 4.2 (Non-Minimal State).
The non-minimal state is:
$X_{\text{ini}}(k)$ $=\mathbf{z}_{\text{ini}}(k)=\text{col}(\mathbf{y}_{\text{ini}}(k),\mathbf{u}_{\text{ini}}(k))=\text{col}(y(k-T_{\text{ini}}),\dots,y(k-1),u(k-T_{\text{ini}}),\dots,u(k-1)),$
with $\mathbf{y}_{\text{ini}}(k)\in\mathbb{R}^{n_{y}T_{\text{ini}}},\mathbf{u}_{\text{ini}}(k)\in\mathbb{R}^{n_{u}T_{\text{ini}}}$. The error state is:
$\tilde{X}_{\text{ini}}(k)$ $=\text{col}(\mathbf{y}_{\text{ini}}(k)-\mathbf{r}_{y},\mathbf{u}_{\text{ini}}(k)-\mathbf{r}_{u}),$
where $\mathbf{r}_{y}=\text{col}(r_{y},\dots,r_{y})$, $\mathbf{r}_{u}=\text{col}(r_{u},\dots,r_{u})$.
Definition 4.3 (Class $\mathcal{K}$ and $\mathcal{KL}$ Functions).
A function $\varphi:\mathbb{R}_{+}\to\mathbb{R}_{+}$ is of class $\mathcal{K}$ if it is continuous, strictly increasing, and $\varphi(0)=0$. It is $\mathcal{K}_{\infty}$ if $\lim_{s\to\infty}\varphi(s)=\infty$. A function $\beta:\mathbb{R}_{+}\times\mathbb{R}_{+}\to\mathbb{R}_{+}$ is $\mathcal{KL}$ if $\beta(\cdot,t)\in\mathcal{K}$ for fixed $t\geq 0$, and $\beta(s,t)\to 0$ as $t\to\infty$ for fixed $s>0$. The identity is $\text{id}(s)=s$.
Definition 4.4 (Filtration).
The filtration $\mathcal{F}_{k}$ is the sigma-algebra generated by the measurements $\{\mathbf{z}_{\text{ini},m}(t)\}_{t=0}^{k}$, representing the information available up to time $k$.
4.1 Noise-Free Case
We analyze the stability of NTDPC in the noise-free case ($\mathbf{\zeta}_{\text{ini}}(k)=0$), tracking references $(r_{y},r_{u})$.
Assumption 4.1 (Positive Definiteness).
There exist $\alpha_{1},\alpha_{2}\in\mathcal{K}_{\infty}$ such that for all $(y,u)\in\mathcal{Y}_{i}\times\mathcal{U}_{i}$, where $\mathcal{Y}_{i}\subseteq\mathbb{R}^{n_{y}}$ and $\mathcal{U}_{i}\subseteq\mathbb{R}^{n_{u}}$ are the projections of $\mathcal{Y}$ and $\mathcal{U}$ onto the $i$-th prediction step:
$\alpha_{1}(\|\text{col}(y-r_{y},u-r_{u})\|)\leq l(y,u)\leq\alpha_{2}(\|\text{col}(y-r_{y},u-r_{u})\|),$
where $l(y,u)=\|y-r_{y}\|_{Q_{i}}^{2}+\|u-r_{u}\|_{R_{i}}^{2}$, and $(r_{y},r_{u})\in\mathcal{Y}_{i}\times\mathcal{U}_{i}$ is a constant admissible equilibrium.
Assumption 4.2 (Terminal Stabilizing Condition).
For any admissible initial state $\mathbf{z}_{\text{ini}}(k)\in\mathcal{U}^{T_{\text{ini}}}\times\mathcal{Y}^{T_{\text{ini}}}$, there exist $\rho\in\mathcal{K}_{\infty}$ with $\rho<\text{id}$, a prediction horizon $N\geq T_{\text{ini}}$, and $\hat{u}(N|k)\in\mathcal{U}_{N}\subseteq\mathbb{R}^{n_{u}}$ such that $\hat{y}(N|k)\in\mathcal{Y}_{N}\subseteq\mathbb{R}^{n_{y}}$:
$l(\hat{y}(N|k),\hat{u}(N|k))\leq(\text{id}-\rho)\circ l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}})).$
Assumption 4.3 (Feasibility).
Problem 4 is feasible for all $k\geq T_{\text{ini}}$ with $\mathbf{\zeta}_{\text{ini}}(k)=0$, and the optimal slack satisfies $\sigma_{y}^{*}(k)=0$. The sets $\mathcal{U},\mathcal{Y}$ contain the equilibrium $(r_{y},r_{u})$.
Theorem 4.4.
Under Assumptions 4.1, 4.2, and 4.3, with $\mathbf{\zeta}_{\text{ini}}(k)=0$ and a constant admissible equilibrium $(r_{y},r_{u})$, the closed-loop system under NTDPC is asymptotically stable, i.e., $y(k)\to r_{y}$, $u(k)\to r_{u}$ as $k\to\infty$.
Proof.
Define the non-minimal state as in Definition 4.2:
$\mathbf{z}_{\text{ini}}(k)=\text{col}(y(k-T_{\text{ini}}),\ldots,y(k-1),u(k-T_{\text{ini}}),\ldots,u(k-1)).$
The error state is $\tilde{X}_{\text{ini}}(k)=\mathbf{z}_{\text{ini}}(k)-\text{col}(\mathbf{r}_{y},\mathbf{r}_{u}),$ where $\mathbf{r}_{y}=\text{col}(r_{y},\ldots,r_{y})$, $\mathbf{r}_{u}=\text{col}(r_{u},\ldots,r_{u})$. Define the auxiliary function as $W(\mathbf{z}_{\text{ini}}(k))=\sum_{i=1}^{T_{\text{ini}}}l(y(k-i),u(k-i)),$ where the stage cost is $l(y(i|k),u(i|k))=\|y(i|k)-r_{y}\|_{Q_{i}}^{2}+\|u(i|k)-r_{u}\|_{R_{i}}^{2},$ with $Q_{i},R_{i}\succ 0$. The optimal cost of Problem 4 is $J^{*}(k)=\sum_{i=0}^{N-1}l(y^{*}(i|k),u^{*}(i|k)),$ since $\sigma_{y}^{*}(k)=0$ in the noise-free case (Assumption 4.3). Consider the storage function [24]:
$V(\mathbf{z}_{\text{ini}}(k))=J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k)).$
To show positive definiteness of $V(\mathbf{z}_{\text{ini}}(k))$, using Assumption 4.1, we have:
$l(y(k-i),u(k-i))\geq\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|),$
so $W(\mathbf{z}_{\text{ini}}(k))\geq\sum_{i=1}^{T_{\text{ini}}}\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|).$ Since $\tilde{X}_{\text{ini}}(k)=\text{col}(y(k-T_{\text{ini}})-r_{y},\ldots,u(k-1)-r_{u})$, there exists $\alpha_{1,W}\in\mathcal{K}_{\infty}$ such that $W(\mathbf{z}_{\text{ini}}(k))\geq\alpha_{1,W}(\|\tilde{X}_{\text{ini}}(k)\|).$ Since $J^{*}(k)\geq 0$, there exists $\alpha_{1,V}\in\mathcal{K}_{\infty}$ such that $V(\mathbf{z}_{\text{ini}}(k))\geq\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|).$ For the upper bound:
$l(y(k-i),u(k-i))\leq\alpha_{2}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|),$
so $W(\mathbf{z}_{\text{ini}}(k))\leq T_{\text{ini}}\alpha_{2}\left(\max_{i=1,\dots,T_{\text{ini}}}\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|\right).$ For $J^{*}(k)$, the predictor output $\hat{\mathbf{y}}_{N}^{*}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini}}(k)+\mathbf{P}_{2}\mathbf{u}_{N}^{*}(k)$ (Lemma 3.5) depends on $\mathbf{z}_{\text{ini}}(k)=\tilde{X}_{\text{ini}}(k)+\text{col}(\mathbf{r}_{y},\mathbf{r}_{u})$. Bounded sets $\mathcal{Y},~{}\mathcal{U}$ (Assumption 4.3) and the predictor’s linearity ensure that $\|y^{*}(i|k)-r_{y}\|$, $\|u^{*}(i|k)-r_{u}\|$ are bounded by $\|\tilde{X}_{\text{ini}}(k)\|$. Thus, there exists $\alpha_{2,J}\in\mathcal{K}_{\infty}$ such that $J^{*}(k)\leq\alpha_{2,J}(\|\tilde{X}_{\text{ini}}(k)\|).$ Combining, there exists $\alpha_{2,V}\in\mathcal{K}_{\infty}$ such that $V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|).$ Hence, $\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|)\leq V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|).$ Next, by defining the supply function as:
$s(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))=l(\hat{y}(N|k),\hat{u}(N|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}})),$
where $\hat{y}(N|k)$, $\hat{u}(N|k)$ satisfy Assumption 4.2. Construct a suboptimal solution at $k+1$:
$\mathbf{u}_{s}(k+1)$ $=\text{col}(u^{*}(1|k),\ldots,u^{*}(N-1|k),\hat{u}(N|k))$
$\mathbf{y}_{s}(k+1)$ $=\text{col}(y^{*}(1|k),\ldots,y^{*}(N-1|k),\hat{y}(N|k)),$ (53)
with $\sigma_{y,s}(k+1)=0$. The suboptimal cost is as $J_{s}(k+1)=\sum_{j=1}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+l(\hat{y}(N|k),\hat{u}(N|k)).$ Since $J^{*}(k)=\sum_{j=0}^{N-1}l(y^{*}(j|k),u^{*}(j|k))$ and $J^{*}(k+1)\leq J_{s}(k+1)$:
$J^{*}(k+1)\leq J^{*}(k)-l(y^{*}(0|k),u^{*}(0|k))+l(\hat{y}(N|k),\hat{u}(N|k))$ (54)
For the auxiliary function $W$:
$W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))=l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))$ (55)
Using (61) and (55), thus:
$V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))=[J^{*}(k+1)-J^{*}(k)]+[W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))]$ (56)
Then, we can rewrite the preceding equality in (56) as:
$V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))$ $\leq[J^{*}(k)-l(y^{*}(0|k),u^{*}(0|k))+l(\hat{y}(N|k),\hat{u}(N|k))]$
$\quad-J^{*}(k)+[l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))]$
$=s(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))$ (57)
By Assumption 4.2, $s(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))\leq-\rho\circ l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}})).$ Since $l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))\geq\alpha_{1}(\|\text{col}(y(k-T_{\text{ini}})-r_{y},u(k-T_{\text{ini}})-r_{u})\|)$, we can rewrite the inequality (57) as:
$V(\mathbf{z}_{\text{ini}}(k+1))-V(\mathbf{z}_{\text{ini}}(k))\leq-\rho\circ\alpha_{1}(\|\text{col}(y(k-T_{\text{ini}})-r_{y},u(k-T_{\text{ini}})-r_{u})\|)\leq-\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|).$
Since $V$ is positive definite and radially unbounded with respect to $\tilde{X}_{\text{ini}}(k)$, and satisfies a strict dissipation inequality, standard Lyapunov results [25] imply $\tilde{X}_{\text{ini}}(k)\to 0$ as $k\to\infty$. Thus, $y(k)\to r_{y}$, $u(k)\to r_{u}$, proving asymptotic stability. Since $V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)$ and $W(\mathbf{z}_{\text{ini}}(k))\geq 0$:
$J^{*}(k)\leq V(\mathbf{z}_{\text{ini}}(k))\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|),$
ensuring the cost is bounded by a $\mathcal{K}_{\infty}$ function of the error state. This completes the proof. ∎
Remark 4.1.
The bound $J^{*}(k)\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)$ satisfies the requirement for a $\mathcal{K}_{\infty}$ cost bound, derived within the storage function analysis.
4.2 Noisy Case: Robustness to Measurement Noise
We analyze the stability of NTDPC under measurement noise ($\mathbf{\zeta}_{\text{ini}}(k)\neq 0$), tracking constant admissible references $(r_{y},r_{u})$. The system is shown to be input-to-state stable (ISS) with respect to $\mathbf{\zeta}_{\text{ini}}(k)$, following the framework in [2, 25].
Assumption 4.5 (Bounded Noise Variance).
There exists $\bar{c}>0$ such that $\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}|\mathcal{F}_{k}]\leq\bar{c}$ for all $k$, where $\mathcal{F}_{k}$ is the filtration up to time $k$.
Assumption 4.6 (Lipschitz Predictor).
There exists $L_{P}>0$ such that:
$\|\mathbf{P}_{1}(\mathbf{z}_{\text{ini},m}(k)-\mathbf{z}_{\text{ini}}(k))\|\leq L_{P}\|\mathbf{\zeta}_{\text{ini}}(k)\|,$
where $L_{P}=\|\mathbf{P}_{1}\|_{2}$ is the spectral norm of $\mathbf{P}_{1}$, and $\mathbf{z}_{\text{ini},m}(k)=\mathbf{z}_{\text{ini}}(k)+\text{col}(0,\mathbf{\zeta}_{\text{ini}}(k))$.
Assumption 4.7 (Terminal Stabilizing Condition - Noisy).
For any admissible initial state $\mathbf{z}_{\text{ini}}(k)\in\mathcal{U}^{T_{\text{ini}}}\times\mathcal{Y}^{T_{\text{ini}}}$, there exist $\rho\in\mathcal{K}_{\infty}$ with $\rho<id$, a prediction horizon $N\geq T_{\text{ini}}$, and $\hat{u}(N|k)\in\mathcal{U}_{N}\subseteq\mathbb{R}^{n_{u}}$ such that $\hat{y}(N|k)\in\mathcal{Y}_{N}\subseteq\mathbb{R}^{n_{y}}$:
$\mathbb{E}[l(\hat{y}(N|k),\hat{u}(N|k))|\mathcal{F}_{k}]\leq(\text{id}-\rho)\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}].$
Assumption 4.8 (Feasibility - Noisy).
Problem 4 is feasible with probability at least $1-\epsilon$, for small $\epsilon>0$. There exists $c_{\sigma}>0$ such that:
$\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}|\mathcal{F}_{k}].$
The sets $\mathcal{U},\mathcal{Y}$ contain the equilibrium $(r_{y},r_{u})$.
Remark 4.2.
Assumption 4.7 is satisfied by selecting $\hat{u}(N|k)=r_{u}$ and leveraging the predictor’s robustness (Lemma 3.5). Assumption 4.8 is supported by the sensitivity index $I_{s}\leq 0.7$ (Fig. 5), ensuring noise separability and feasibility.
Theorem 4.9.
Under Assumptions 4.1, 4.5, 4.6, 4.7, and 4.8, with a constant admissible equilibrium $(r_{y},r_{u})$, NTDPC is input-to-state stable (ISS) with respect to $\mathbf{\zeta}_{\text{ini}}(k)$. There exist $\beta\in\mathcal{KL}$, $\gamma\in\mathcal{K}_{\infty}$ such that:
$\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|]\leq\beta(\|\tilde{X}_{\text{ini}}(0)\|,k)+\gamma\left(\sup_{0\leq t<k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}\right).$
Proof.
Define the non-minimal state as in Definition 4.2:
$\mathbf{z}_{\text{ini}}(k)=\text{col}(y(k-T_{\text{ini}}),\ldots,y(k-1),u(k-T_{\text{ini}}),\ldots,u(k-1)),$
and the error state $\tilde{X}_{\text{ini}}(k)=\mathbf{z}_{\text{ini}}(k)-\text{col}(\mathbf{r}_{y},\mathbf{r}_{u}),$ where $\mathbf{r}_{y}=\text{col}(r_{y},\ldots,r_{y})$, $\mathbf{r}_{u}=\text{col}(r_{u},\ldots,r_{u})$. The auxiliary function is $W(\mathbf{z}_{\text{ini}}(k))=\sum_{i=1}^{T_{\text{ini}}}l(y(k-i),u(k-i)),$ with stage cost $l(y(i|k),u(i|k))=\|y(i|k)-r_{y}\|_{Q_{i}}^{2}+\|u(i|k)-r_{u}\|_{R_{i}}^{2},$ and $Q_{i},R_{i}\succ 0$. The optimal cost of Problem 4 is $J^{*}(k)=\sum_{i=0}^{N-1}l(y^{*}(i|k),u^{*}(i|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}.$ Define the stochastic Lyapunov function:
$V(k)=\mathbb{E}[J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}].$
In order to study positive definiteness of $V(k)$, under Assumption 4.1, we have $l(y(k-i),u(k-i))\geq\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|),$ so $W(\mathbf{z}_{\text{ini}}(k))\geq\sum_{i=1}^{T_{\text{ini}}}\alpha_{1}(\|\text{col}(y(k-i)-r_{y},u(k-i)-r_{u})\|)\geq\alpha_{1,W}(\|\tilde{X}_{\text{ini}}(k)\|),$ for some $\alpha_{1,W}\in\mathcal{K}_{\infty}$. For $J^{*}(k)$, the predictor output $\hat{\mathbf{y}}_{N}^{*}(k)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k)+\mathbf{P}_{2}\mathbf{u}_{N}^{*}(k)+\sigma_{y}^{*}(k)$ and bounded sets $\mathcal{Y},\mathcal{U}$ (Assumption 4.8) imply that $\|y^{*}(i|k)-r_{y}\|$, $\|u^{*}(i|k)-r_{u}\|$ are bounded by $\|\tilde{X}_{\text{ini}}(k)\|$ and noise effects. By Assumption 4.8:
$\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k)\|^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\bar{c}.$
Thus, there exists $\alpha_{2,J}\in\mathcal{K}_{\infty}$ such that:
$\mathbb{E}[J^{*}(k)|\mathcal{F}_{k}]\leq\alpha_{2,J}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}.$
Combining, there exist $\alpha_{1,V},\alpha_{2,V}\in\mathcal{K}_{\infty}$ such that:
$\alpha_{1,V}(\|\tilde{X}_{\text{ini}}(k)\|)\leq V(k)\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}.$
As dissipation inequality, we construct a suboptimal solution at $k+1$ like (53) with $\sigma_{y,s}(k+1)=\mathbf{P}_{1}\mathbf{z}_{\text{ini},m}(k+1)+\mathbf{P}_{2}\mathbf{u}_{s}(k+1)-\mathbf{y}_{s}(k+1)$, where $\hat{u}(N|k),\hat{y}(N|k)$ satisfy Assumption 4.7. The suboptimal cost is:
$J_{s}(k+1)=\sum_{j=1}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+l(\hat{y}(N|k),\hat{u}(N|k))+\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}.$
By computing $\mathbb{E}[J_{s}(k+1)|\mathcal{F}_{k}]$ as follows:
$\mathbb{E}[J_{s}(k+1)|\mathcal{F}_{k}]=\mathbb{E}\left[\sum_{j=1}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+l(\hat{y}(N|k),\hat{u}(N|k))+\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}\Big{|}\mathcal{F}_{k}\right]$ (58)
Since:
$\mathbb{E}[J^{*}(k)|\mathcal{F}_{k}]=\mathbb{E}\left[\sum_{j=0}^{N-1}l(y^{*}(j|k),u^{*}(j|k))+\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}\Big{|}\mathcal{F}_{k}\right]$ (59)
Then, by subtracting (58) and (59), we have:
$\mathbb{E}[J_{s}(k+1)-J^{*}(k)|\mathcal{F}_{k}]$ $=-\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]-\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]$
$\quad+\mathbb{E}[l(\hat{y}(N|k),\hat{u}(N|k))|\mathcal{F}_{k}]+\mathbb{E}[\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]$
By Assumption 4.7:
$\mathbb{E}[l(\hat{y}(N|k),\hat{u}(N|k))|\mathcal{F}_{k}]\leq(id-\rho)\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}].$ (60)
For the slack terms, by Assumption 4.6:
$\|\sigma_{y,s}(k+1)\|\leq\|\mathbf{P}_{1}(\mathbf{z}_{\text{ini},m}(k+1)-\mathbf{z}_{\text{ini}}(k+1))\|\leq L_{P}\|\mathbf{\zeta}_{\text{ini}}(k+1)\|$ (61)
so, by taking the expectation from both side of (61):
$\mathbb{E}[\|\sigma_{y,s}(k+1)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq\lambda_{\max}(\Lambda_{y})L_{P}^{2}\mathbb{E}[\|\mathbf{\zeta}_{\text{ini}}(k+1)\|^{2}|\mathcal{F}_{k}]\leq\lambda_{\max}(\Lambda_{y})L_{P}^{2}\bar{c}.$
Similarly, $\mathbb{E}[\|\sigma_{y}^{*}(k)\|_{\Lambda_{y}}^{2}|\mathcal{F}_{k}]\leq c_{\sigma}\bar{c}$. Let $c_{\sigma,\text{max}}=\max(c_{\sigma},\lambda_{\max}(\Lambda_{y})L_{P}^{2})$. Then:
$\mathbb{E}[J_{s}(k+1)-J^{*}(k)|\mathcal{F}_{k}]\leq-\rho\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}]+2c_{\sigma,\text{max}}\bar{c}.$ (62)
And for $W$:
$W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))=l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))$ (63)
By taking the expectation from both side of (63), thus:
$\mathbb{E}[W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}]=\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))-l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}].$ (64)
By combining (62) and (64):
$V(k+1)-V(k)\leq\mathbb{E}[J_{s}(k+1)-J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k+1))-W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}].$ (65)
By substituting (60) in (65), yields:
$V(k+1)-V(k)\leq-\rho\circ\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}]+\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]+2c_{\sigma,\text{max}}\bar{c}.$
By Assumption 4.1, there exists $\alpha_{4}\in\mathcal{K}_{\infty}$ such that:
$\mathbb{E}[l(y(k-T_{\text{ini}}),u(k-T_{\text{ini}}))|\mathcal{F}_{k}]\geq\alpha_{4}(\mathbb{E}[\|\text{col}(y(k-T_{\text{ini}})-r_{y},u(k-T_{\text{ini}})-r_{u})\|])$ (66)
Since $\tilde{X}_{\text{ini}}(k)$ includes these terms and using (66), there exists $\alpha_{4,V}\in\mathcal{K}_{\infty}$ such that:
$V(k+1)-V(k)\leq-\rho\circ\alpha_{4,V}(\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|])+\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]+2c_{\sigma,\text{max}}\bar{c}.$
The term $\mathbb{E}[l(y^{*}(0|k),u^{*}(0|k))|\mathcal{F}_{k}]$ is bounded by the predictor’s output and noise effects, contributing to the ISS gain.
To show ISS property, By Assumption 4.5, $\bar{c}\leq c_{\zeta}\sup_{t\leq k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}^{2}$, the dissipation inequality:
$V(k+1)\leq V(k)-\rho\circ\alpha_{4,V}(\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|])+\alpha_{5}(\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|])+c_{\zeta,\text{max}}\sup_{t\leq k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}^{2},$
where $\alpha_{5}\in\mathcal{K}_{\infty}$ bounds the control cost, and $c_{\zeta,\text{max}}=2c_{\sigma,\text{max}}c_{\zeta}$. Applying stochastic ISS results [25], there exist $\beta\in\mathcal{KL}$, $\gamma\in\mathcal{K}_{\infty}$ such that:
$\mathbb{E}[\|\tilde{X}_{\text{ini}}(k)\|]\leq\beta(\|\tilde{X}_{\text{ini}}(0)\|,k)+\gamma\left(\sup_{0\leq t<k}\|\mathbf{\zeta}_{\text{ini}}(t)\|_{\infty}\right).$
Since $V(k)\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c}$ and $\mathbb{E}[W(\mathbf{z}_{\text{ini}}(k))|\mathcal{F}_{k}]\geq 0$:
$\mathbb{E}[J^{*}(k)|\mathcal{F}_{k}]\leq\alpha_{2,V}(\|\tilde{X}_{\text{ini}}(k)\|)+c_{\sigma}\bar{c},$
ensuring the cost is bounded by a $\mathcal{K}_{\infty}$ function of the error state plus a noise-dependent term. Hence, the proof is completed. ∎
Remark 4.3.
The ISS property ensures bounded tracking errors under noise, as validated in Fig. 2, where NTDPC maintains stable tracking of $(r_{y},r_{u})$ with $\sigma_{\zeta}^{2}=0.25I$. The sensitivity index $I_{s}$ (Fig. 5) supports Assumption 4.8, aiding horizon selection.
5 Numerical Simulations
Consider the discrete-time linear model system we aim to control, characterized by the following matrices [26], Section 6.4:
$\begin{split}&A=\begin{bmatrix}0.9997&0.0038&-0.0001&-0.0322\\ -0.0056&0.9648&0.7446&0.0001\\ 0.0020&-0.0097&0.9543&-0.0000\\ 0.0001&-0.0005&0.0978&1.0000\end{bmatrix},\quad\\ &B=\begin{bmatrix}0.0010&0.1000\\ -0.0615&0.0183\\ -0.1133&0.0586\\ -0.0057&0.0029\end{bmatrix},C=\begin{bmatrix}1.0000&0&0&0\\ 0&-1.0000&0&7.7400\end{bmatrix}\end{split}$ (67)
These matrices are obtained by discretizing the dynamics of the longitudinal motion of a Boeing 747 using a zero-order hold with a sampling time $T_{s}=0.1\text{ s}$. The system has two inputs: the throttle $u_{1}$ and the elevator angle $u_{2}$, and two outputs: the aircraft’s longitudinal velocity and climb rate.
The control objective is to increase the longitudinal velocity ($y_{1}(t)$) from 0 ft/s to 10 ft/s while maintaining the climb rate ($y_{2}(t)$) at zero. This corresponds to a reference output $r_{y}=\begin{bmatrix}10&0\end{bmatrix}^{T}$. The inputs and outputs are constrained as follows:
$\begin{split}&U:=\left\{u\in\mathbb{R}^{2}:\begin{bmatrix}-20\\ -20\end{bmatrix}\leq u\leq\begin{bmatrix}20\\ 20\end{bmatrix}\right\},\quad\\ &Y:=\left\{y\in\mathbb{R}^{2}:\begin{bmatrix}-25\\ -15\end{bmatrix}\leq y\leq\begin{bmatrix}25\\ 15\end{bmatrix}\right\}.\end{split}$ (68)
In this simulation, we set $T_{\text{ini}}=N=20$, data length for Hankel matrices $M=2500$.
5.1 Noise-free case
The Fig.1 illustrates the performance of NTDPC in the noise-free case ($\mathbf{\zeta}_{\text{ini}}(k)=0$) when tracking constant admissible references ($(r_{y},r_{u})$), as analyzed in Section 5.1. The results are consistent with the asymptotic stability proven in Theorem 4.4, where $y(k)\to r_{y}$ and $u(k)\to r_{u}$ as $k\to\infty$ under Assumptions 4.1, 4.2, and 4.3. The top plot displays the primary output $y_{1}(k)$ for NTDPC (blue), SPC (red), and SMMPC (green) over time, with the dashed line representing the reference $r_{y}$; all methods converge to $r_{y}$, but NTDPC exhibits the fastest transient response and minimal overshoot, reflecting the effectiveness of the data-driven predictor and terminal condition. The second plot shows a secondary output $y_{2}(k)$, with similar convergence behavior, where NTDPC (blue) stabilizes more quickly than SPC (red) and SMMPC (green), supporting the robustness of the NTDPC framework to initial conditions encoded in $\mathbf{z}_{\text{ini}}(k)$. The third plot presents the primary input $u_{1}(k)$, where NTDPC (blue) adjusts rapidly from an initial value, stabilizing near $r_{u}$ (dashed line) with less oscillation compared to SPC (red) and SMMPC (green), consistent with the bounded input constraints in $\mathcal{U}$. Finally, the bottom plot depicts a secondary input $u_{2}(k)$, showing NTDPC (blue) achieving a steady state closer to $r_{u}$ with reduced variability, further validating the stability properties derived from the Lyapunov function $V(\mathbf{z}_{\text{ini}}(k))$. Overall, the convergence of outputs and inputs to their respective references demonstrates the asymptotic stability of NTDPC, as predicted by the storage function analysis $V(\mathbf{z}_{\text{ini}}(k))=J^{*}(k)+W(\mathbf{z}_{\text{ini}}(k))$, and the superior performance of NTDPC over SPC and SMMPC highlights the advantage of the noise-tolerant predictor and optimization framework, particularly in the absence of noise.
Figure 1: Output trajectories for noise-free case
5.2 Noisy case
In Fig. 2, the simulation results demonstrate the performance of NTDPC in comparison to SPC and SMMPC under the noise effect, $\sigma_{\zeta}^{2}=0.25I$. While SMMPC exhibits aggressive and unstable control actions leading to poor regulation of outputs, NTDPC and SPC show reasonable results, despite an initial overshoot and undershoot, ultimately achieves stable tracking of the reference for both outputs with damped oscillations and less aggressive control inputs that settle over time. This balanced response in both outputs and the settling behavior of the control inputs highlight the effectiveness and stability of the NTDPC and SPC strategies compared to the oscillatory and poorly regulated responses observed with SMMPC.
Figure 2: Output trajectories over 50 Monte Carlo runs for noisy case
To quantify controller performance, the objective function is evaluated and plotted in Fig. 3 which reveals the superiority of NTDPC and SPC compared to SMMPC. It is worthy to note that there is a slight difference between NTDPC and SPC throughout the entire simulation, highlighting the effectiveness of the designed NTDPC. In summary, the figures provide a comprehensive comparison of NTDPC, SPC and SMMPC control strategies, highlighting their performance in terms of output and input trajectories and their respective variability. This comparison aids in understanding the strengths and weaknesses of each method in maintaining desired control objectives amidst system noise and constraints.
Figure 3: Performance index over 50 Monte Carlo runs for noisy case
5.3 Sensitivity Evaluation
Increasing horizon lengths improves estimation quality but increases computational cost and adaptation time in time-varying systems. We propose a sensitivity index to guide parameter selection under various noise conditions. When using noisy data, matrix $\Sigma_{2}$ in (18) becomes non-zero. Ideally, $\Sigma_{2}$’s singular values should be noise-related and significantly smaller than those of $\Sigma_{1}$. In practice, both matrices contain signal and noise components. Figure 4 shows singular value patterns for $\Sigma_{1}$ and $\Sigma_{2}$ with $\sigma_{\zeta}^{2}=0.01I$ and $T_{\text{ini}}=N=15$, with a vertical line separating the two sets.
Figure 4: Singular value distribution for $T_{\text{ini}}=N=15$ and $\sigma_{\zeta}^{2}=0.01I$
To quantify signal-noise separability, we propose:
$I_{s}=\frac{\sigma_{\text{max}}^{2}(\Sigma_{2})}{\sigma_{\text{min}}^{2}(\Sigma_{1})}$ (69)
Lower $I_{s}$ values indicate better signal-noise separability. Figure 5 shows $I_{s}$ behavior as $T_{\text{ini}}$ increases from 10 to 50 with noise variances from 0.01 to 0.32. Empirically, algorithm convergence requires $I_{s}\lesssim 0.7$. This confirms higher horizons are needed for higher noise variances and guides horizon selection. Each point represents the mean from 50 Monte Carlo simulations, with shading indicating one standard deviation. Since these analyses can be performed before controller implementation (or offline for LTI systems), this approach provides practical guidance for horizon selection.
Figure 5: Sensitivity index for different noises and $T_{\text{ini}}$ values
5.4 Computational Complexity
In this section, we will evaluate computational complexity of the developed NTDPC compared to SPC. In the context of a time-varying system, the prediction model needs to be updated at each sample time. Comparing the computational effort for updating this model between SPC and the NTDPC, focusing solely on the more costly calculations of pseudoinverse and SVD, reveals the following for a simplified SISO system with input/output horizons $T_{\text{ini}}=N=T_{h}$ (using $T_{h}$ to avoid confusion with data length $T$) and a data window length $M$:
For SPC, estimating the prediction matrices involves computing the pseudoinverse of the Hankel matrix $\begin{bmatrix}\mathbf{U}_{p}&\mathbf{Y}_{p}&\mathbf{U}_{f}\end{bmatrix}^{T}$ (or its non-transposed version depending on formulation), which has dimensions $3T_{h}\times M$. Assuming $M\gg 3T_{h}$, the computational complexity of the pseudoinverse is approximately $\mathcal{O}((3T_{h})^{2}M)=\mathcal{O}(9T_{h}^{2}M)$.
For the NTDPC method, the matrix estimation per sample time primarily involves two SVD computations: one for the matrix $Z_{p}$ (dimensions $2T_{h}\times M$) and another for the matrix $Z_{f}V_{2}$ (dimensions approximately $2T_{h}\times(M-T_{h}-n)$, where $n$ is the system order, here simplified to $2T_{h}\times(M-T_{h})$). The computational complexity of the SVDs are approximately $\mathcal{O}((2T_{h})^{2}M)$ and $\mathcal{O}((2T_{h})^{2}(M-T_{h}-n))$, respectively, assuming the number of rows is less than the number of columns in both cases ($2T_{h}<M$ and $2T_{h}<M-T_{h}-n$). Summing these, the total dominant SVD cost for NTDPC is approximately $\mathcal{O}(4T_{h}^{2}M+4T_{h}^{2}(M-T_{h}-n))\approx\mathcal{O}(8T_{h}^{2}M)$ for large $M$.
This comparison of the dominant factorization/inversion costs highlights that NTDPC offers a reduction in computational effort per sample time with a lower constant factor (approximately 8 compared to 9) in the $\mathcal{O}(T_{h}^{2}M)$ complexity. Furthermore, similar to other hybrid methods the NTDPC works with reduced order Hankel matrices which in turn reduces the memory needed for the online matrix estimation step, contributing to improved overall efficiency for time-varying systems.
6 Conclusion
This paper presents a comprehensive analysis of the Noise-Tolerant Data-Driven Predictive Control (NTDPC) framework, addressing the critical challenge of measurement noise in hybrid data-driven predictive control while achieving robust tracking of non-zero references. By leveraging singular value decomposition (SVD), NTDPC effectively separates system dynamics from noise within reduced-order Hankel matrices, enabling accurate trajectory predictions with shorter data horizons and reduced computational complexity. The stability analysis demonstrates that NTDPC achieves asymptotic stability in the noise-free case, ensuring precise tracking of the reference equilibrium. In the presence of measurement noise, NTDPC exhibits input-to-state stability (ISS), guaranteeing robust performance with bounded tracking errors proportional to the noise magnitude.
Numerical simulations on a flight control benchmark validate the theoretical findings, showcasing NTDPC’s superior performance compared to SMMPC and SPC. The output trajectories demonstrate stable tracking with damped oscillations and less aggressive control inputs, while the performance index highlights NTDPC’s effectiveness in minimizing tracking errors. The proposed sensitivity index, $I_{s}$, provides a practical tool for selecting prediction horizons under varying noise conditions, with empirical results indicating that $I_{s}\lesssim 0.7$ ensures algorithm convergence. This index, computed offline for linear time-invariant systems, enhances the applicability of NTDPC in real-world scenarios. Furthermore, the computational complexity analysis reveals that NTDPC reduces the dominant factorization costs by approximately 11% compared to SPC, with lower memory requirements due to reduced-order Hankel matrices, making it well-suited for time-varying systems.